OpenStudy (anonymous):

Jonah was keeping track of how many cars drove by every hour. He realized that the number of cars were in a sequence that followed the pattern n + 5. How many cars drove by on the 15th hour? 75 10 3 20

OpenStudy (anonymous):

please help?

OpenStudy (anonymous):

In the first hour, n =1 and n + 5 = 6. In the second hour, n=2 and n + 5 = 7. 3rd hour, n = 3 and n + 5 = 8. This will establish a pattern and you can see what n and n + 5 will be for the 15th hour now.

OpenStudy (anonymous):

so wat is the answer????

OpenStudy (anonymous):

HELPPPPPPP!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Show me some of your work. You are supposed to be the one doing the work and we tutors are only supposed to help and guide.

OpenStudy (anonymous):

this is multiple question!

OpenStudy (anonymous):

it is 20

OpenStudy (anonymous):

Good! See, you were able to do it after all and just needed a little push. You got it! Good job!

OpenStudy (anonymous):

i have another question!

OpenStudy (anonymous):

pwease help

OpenStudy (anonymous):

]Which expression results from applying the Identity Property of Multiplication to s + 9s + 7s + 11 to get an equivalent expression?

OpenStudy (anonymous):

1s + 9s + 7s + 11 s + 16s + 11 s + s(9 + 7) + 11 s + 9s + 7(s+11)

OpenStudy (anonymous):

The identity property of multiplication says that any number multiplied by 1 is the same number. From among your options, there is only one equation exhibiting that. With this explanation you should be able to find it.

OpenStudy (anonymous):

1s+9s+7s+11?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

u there????

OpenStudy (anonymous):

u there HELLO!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

You got the answer! Another good job.